JAN342gg:

Alex is six years older than his sister, Emma. The sum of their ages is 32. How old is Alex?

JAN342gg:

26

Eiwoh2:

Correct. So I started this wrong. The way to do it is to assume Alex is X years old and Emma is Y years old. And X + Y = 32, since 32’s the sum.

Eiwoh2:

The key to this is substitution. X - 6 = Y X + Y = 32. Do you know what to do with these equations?

silvernight269:

@JAN342gg